### v2

# Large Model Strategic Thinking, Small Model Efficiency: Transferring Theory of Mind in Large Language Models

Abstract
As the performance of larger, newer Large Language Models continues to improve for strategic Theory of Mind (ToM) tasks, the demand for these state of the art models increases commensurately. However, their deployment is costly both in terms of processing power and time. In this paper, we investigate the feasibility of creating smaller, highly-performing specialized algorithms by way of fine-tuning. To do this, we present a large pre-trained model with 20 unique scenarios that combine a social context with a social dilemma, recording its answers, and using them for Q&A fine-tuning on a smaller model of the same family. Our focus is on in-context game-theoretic decision-making, the same domain within which human interaction occurs and that requires both a theory of mind (or a semblance thereof) and an understanding of social dynamics. We find that the fine-tuned smaller language model exhibited significant performance closer to that of its larger relative, and that their improvements extended in areas and contexts beyond the ones provided in the training examples. On average for all games, through fine-tuning, the smaller model showed a 46% improvement in aligning with the behavior of the larger model, with 100% representing complete alignment. This suggests that our pipeline represents an efficient method to transmit some form of theory of mind to smaller models, creating improved and cheaply deployable algorithms in the process. Despite their simplicity and their associated shortcomings and limitations, our findings represent a stepping stone in the pursuit and training of specialized models for strategic and social decision making.
1 Introduction
The concept of Theory of Mind (ToM) refers to the ability to impute mental states such as beliefs, intentions or desires to the self and others [1]. As such, ToM has been extensively studied among humans [2] and other animals [3] , but it has recently received a surge of attention with the release of Large Language Models (LLMs) [4, 5, 6, 7]. The latter are sophisticated neural networks based on the Transformers technology, and have displayed impressive capabilities in fields ranging from medicine to finance [8, 9, 10, 11]. Findings are rarely unequivocal, and it is not unheard of that follow-up papers dispute or outright reject the findings of a previous publication [12, 13]. Investigation into the ToM capabilities of these powerful algorithms was naturally prompted by both their excellent performance and their potential for human-AI interaction, however, much like for the rest of LLM research, evidence is still mixed on the subject.
Whereas the usage of a more modern or larger model does not necessarily translate into better performance for ToM tasks [14], there is consistent evidence that newer models with more parameters are better strategic decision makers [15]. While encouraging, this finding also implies that the rising demand for algorithmic advisors [16, 17] would preferentially be served by larger, more computationally demanding and polluting algorithms [18]. Furthermore, it should be stated that within the larger umbrella of ToM tasks, strategic decision making occupies a very delicate niche. Indeed, strategic interaction among humans takes place mostly through natural language [19, 20], the same type of output produced by LLMs, who could therefore masquerade as humans if sufficiently sophisticated.The peculiar nature of this subset of tasks thus raises concerns not just about performance, but also about alignment: a very humanlike and highly performing algorithm from a ToM perspective could still be able to act in a very antisocial manner [21].
One potential way to address this dual dilemma is to turn to the method of fine-tuning. Indeed, since Large Language Models are generalists by design, their usefulness in specific domains of knowledge is not as pronounced as it could be, and neither is it indicative of said models’ general and social intelligence. Through fine-tuning, a model can be partially retrained to improve its performance with respect to a certain dataset or some areas of expertise. This technique is particularly promising for those tasks in which rote memorization is essential for success, as evidenced by findings in medical sciences, and could thus prove to be an avenue to create domain experts "in silico" [22, 23, 24]. On the other hand, its usefulness in open-ended tasks that require strategic sophistication is still unclear: in some cases, not even fine-tuning is sufficient to make up for the shortcomings of state-of-the-art LLMs when it comes to the theory of mind reasoning [25]. The problem is further compounded by the absence of datasets that could be used to fine-tune an LLM with the purpose of making it more strategic without compromising alignment, as well as a lack of clear guidelines for what this dataset should include, or how it should be assembled.
Recent research suggests that these algorithms can at least navigate simple economic and game theoretic scenarios, [26, 27, 28, 29], but their choices of action tend to be anchored to broader contextual cues [30]. Said cues are occasionally extended by the model itself into fully-fledged hypothetical scenarios, suggesting the importance of context in both human and synthetic decision-making. Contextual cues can be thought of as the motivation for why humans in the real world would make decisions that depart from the predictions at the intersection of game theory and rational choice theory, and the algorithms’ responsiveness to them suggests that these machines might already possess some level of social alignment. In particular, there is evidence that Llama2-70b aptly navigates the nuances of the intersection between strategic decision-making and contextual awareness [30].
In light of these findings, we propose a fine-tuning procedure to align the behavior of smaller, more agile algorithms with that of their larger counterpart belonging to the same family. By querying LLaMa2-70b, we obtain a dataset containing motivated responses to four different social dilemmas framed within five different types of social interaction. This dataset is then used to fine-tune LLaMa2-7b, and the performance of the newly fine-tuned model is compared to that of both a pre-trained LLaMa2-70b and a pre-trained LLaMa2-7b. Surprisingly, we find that the smaller fine-tuned model behaves more closely to its larger per-trained counterpart not only for the scenarios upon which it has been trained, but also for new and unknown scenarios that were not present in the training data.
These promising findings indicate that a version of the theory of mind can imparted to smaller models, making them a viable, less computationally demanding alternative to their larger siblings.
The rest of the paper is organized as follows. Section 2 introduces the relevant literature, while Section 3 offers an overview of the methods. Section 3 presents the results, while Section 4 concludes.
2 Literature Review
Within the broad and fast-developing literature surrounding Large Language Models, our research lies at the intersection of two broader subjects: model empowerment through fine-tuning, and model evaluation in the field of strategic reasoning.
For what concerns model empowerment, several interesting examples come from just as many subjects [31, 32, 33, 34, 35, 36], with the overall consensus being that fine-tuning can dramatically improve performance in specific areas of knowledge. More interestingly, literature is developing on the process of training smaller models using larger models [37, 38, 39], either via fine-tuning or through direct intervention. Our main contribution in this area is in the usage of a completely synthetic dataset for fine-tuning purposes on a domain area that requires strategic thought.
What characterizes strategic interaction is a necessity to anticipate the moves of a hypothetical co-player - a challenge far different from simply identifying the correct answer to a test. Experiments with LLMs in place of humans tend to yield results that are inconsistent with rational choice theory, behavioral economics, or both [40, 41, 42]. As such, it is currently debated how and where these models should be evaluated [43, 15]. One increasingly popular methodology involves creating LLM-empowered agents and having them compete versus one another or versus humans, occasionally endowing them with humanlike traits and personality [44, 45, 46]. Unfortunately, this approach is simultaneously too narrow to capture variation in human behavior, and too wide to be generally applicable like game theory is. Nevertheless, there is promise in the usage of these algorithms to mimic human thoughts and behavior [47, 48, 49, 50, 51]. What is more, fine-tuning has been shown to improve the performance of large language models even over tasks that require human-like judgment and social skills [52].
3 Methods
For the purpose of our analysis, we use models belonging to the LLaMa-2 family. Although they no longer represent the state-of-the-art, LLaMA-2 models have displayed the least erratic behavior during our investigation and have proven to be much easier to fine-tune. We elaborate on our decision not to use LLaMa-3 in Section E of SI. Of the three available configurations, we select the 70 billion parameters and the 7 billion parameters models to act as the reference large and small pre-trained model, respectively. We load all algorithms on Northeastern University’s High Performance Cluster (HPC), and we access them via the HuggingFace pipeline. We furthermore make use of the langchain package for Python in order to customize system prompts and inputs. Our experiments are conducted on memoryless algorithms with high temperature, in order to treat each instance and question-answer pair as distinct, and to capture the widest possible range in variation.
We distribute a contextual frame of reference upon which to build and elaborate a strategic response to the LLM through the use of system prompts. System prompts serve as an optimal medium for instilling a defined role and personality in LLMs, as they are not processed as typical user inputs but rather function as default settings. This approach is crucial for our objective of grounding LLM decisions within a realistic context. Leveraging system prompts in this manner enhances the ability to align the responses of LLMs with real-world scenarios while not neglecting strategic considerations.
Furthermore, in order to better establish a connection between existing literature and our studies, we adopt the same contextual prompts presented in [30]. These are: a summit between leaders from two nations ("IR"), a meeting between CEOs of two distinct firms ("biz"), a conference where two industry leaders from different companies commit to environmental regulations ("environment"), a discussion between team members vying for the same promotion ("team"), and a conversation between friends aiming for a compromise ("friendsharing"). To these five "canonical" contexts, we add three more: ("sports"), ("ventcap") and ("roomsharing"). We call "scenario" the combination of a context and a game; 20 scenarios are used for fine-tuning purposes containing every possible combination of social dilemma and canonical context. 20 more out-of-sample scenarios are generated combining the five canonical contexts with differently parametrized social dilemmas (with parametrization constant within dilemmas). Additionally, 12 more out-of-sample scenarios are generated combining our three additional contexts with the four social dilemmas, whose parametrization is identical to training parametrization. Finally, we run one more scenario which is geared towards multi-player games and which is entirely out of sample (both in game and in context).
When testing the performance of pre-trained and fine-tuned models on within-sample scenarios, we run 300 initializations and record the response given by the LLM. Prompts are kept constant across LLMs. For LLaMa2-70b, we explicitly ask the algorithm not only to give us its choice of action, but also to include a motivation for it. Although the motivations provided by the LLM might include imprecise reasoning and are unlikely to reflect its actual decision making process, we still chose to include them in the dataset. Given the existing evidence that its final choices should be on average "correct", the accompanying elaborations should reflect and corroborate that with an explanation grounded in natural language. The resulting triplets of system prompt, game prompt and AI response are then collated within a dataset which is used to fine-tune the smaller LLaMa2-7b model. Additionally, as a robustness check, we ask the larger model to output only its choice of action (without motivation) in order to test for stability of preferences. Once the fine-tuning has been performed, we query the fine-tuned model both on the scenarios it has been trained on, and on the new out-of-sample ones using 300 initializations per scenario.
Our game theoretic considerations and modeling choices are detailed in SI B.II. A description of the fine-tuning procedure followed can be found in SI B.III. Prompts used can be found in SI A.
4 Results
Figure 1 gives an overview of the methods we employ in order to generate, collect, and then analyze our results.
Figure 1: Overview of the methods employed in this paper. We pair all games and scenarios to generate 20 unique combinations, which form the backbone of our dataset. We then submit each combination to each model, and obtain 300 observations per combination. For LLaMa2-70b, we ask for an answer and a motivation; we ask the other models only for their answers. We use the answers coming from LLaMa2-70b to perform LORA on a small, pre-trained LLaMa2-7b. The fine-tuned model is then again queried like the pre-trained model, and once that is done, we collect all data and measure the impact of fine-tuning on preferences.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: Preliminary investigation of differences in responses between LLaMa-7b and LLaMa-70b, grouped by context and game. Clockwise, from the left: propensity to cooperate in LLaMa2-7b grouped by context; propensity to cooperate in LLaMa2-7b grouped by game; propensity to cooperate in LLaMa2-70b grouped by context; propensity to cooperate in LLaMa2-70b grouped by games. Notably, LLaMa2-7b is almost entirely indifferent to context and game and displays a remarkable bias for choosing cooperation.
Before initiating the process of fine-tuning, we conducted an exploratory data analysis to capture the differences in behavior between LLaMa2-7B and LLaMa2-70b. Our results, displayed in figure 2, paint a striking figure: the smallest model in the family is ignorant of both context and game, acts almost randomly, and has a bias for choosing cooperation. Nowhere is this more visible than in the case of playing Prisoner’s Delight under the friendsharing context. Not only does Prisoner’s Delight only admit cooperation from the perspective of a perfectly rational player, but the social environment in which the interaction takes place according to the prompt fed to the machine envisions a scenario in which friends are interacting. In spite of this, LLaMa2-7b chooses to cooperate the least when faced with this combination of game and context when compared to the same game with different contexts. More in general, against all expectations, LLaMa2-7b is cooperating a lot less when playing under the friendsharing context as opposed to nearly all other contexts. The results obtained from LLaMa-70b, on the other hand, are perfectly in line with the existing literature, vindicating our choice to use this LLM as a guide through the intricate interplay existing between social context and strategic play.
4.1 Within Sample Results
Figure 3: Results for LLaMa2-70b pre-trained, LLaMa2-7b pre-trained, and LLaMa2-7b fine-tuned grouped by game. Displayed on the $y$ axis is the propensity to cooperate over the span of 300 runs in a game when given a context. Within each game, and for each context, we observe that the propensity to cooperate with the fine-tuned model almost always lies in between that of the pre-trained model of the same size and the pre-trained model of the largest size. Only four notable exceptions can be found. When playing Harmony under the team context and when playing Snowdrift under the friendsharing context, the fine-tuned model overcorrects and cooperates less than the larger pre-trained model. When playing Prisoner’s Dilemma, both in the friendsharing and in the team context, the fine-tuned model displays no evidence of learning from the larger pre-trained model and instead persists or exacerbates the behavior of the smaller pre-trained model.
Upon inspection of the results shown in Figure 3, it appears immediately evident that the smaller model becomes closer in behavior to the larger model after fine-tuning. This is, however, not universally the case: the fine-tuned model becomes less cooperative in the "friendsharing" scenario when playing Prisoner’s Dilemma and Snowdrift, below both the pre-trained and the fine-tuned model; additionally, when playing Prisoner’s Delight under the "team" context, we observe a decrease in cooperation that exceeds the rate of defection of the larger model. Moreover, we notice aberrant behavior when the fine-tuned LLaMa-7b plays the Prisoner’s Dilemma under the "team" context. In these scenarios, the fine-tuned algorithm exacerbates the behavior of its original, pre-trained counterpart and increases its rate of defection. It is worth pointing out that in two of the four irregular cases observed, the change in behavior of the fine-tuned model vis-a-vis the pre-trained one is in the direction of the larger model, or, in other words, the model fine-tuned to the instructions coming from a larger model follows the general trend set by said larger model, even when this results in a puzzling overcorrection.
Figure 4: Improvement of the small fine-tuned model over the 20 in-sample scenarios.
In order to properly quantify the extent to which the fine-tuned small model approximates the behavior choices of its larger, pre-trained counterpart, we define "improvement" in the following terms:
$\mathbb{I}:=1-\frac{70b-7b_{\text{finetuned}}}{70b-7b_{\text{org}}}\times 100$ (1)
where $7b_{\text{org}}$ refers to the count of cooperation choices in the original version of Llama2-7b model, $7b_{\text{finetuned}}$ refers to the same quantity for the fine-tuned version of the same model, and 70b captures the count of cooperative choices for the large Llama2-70b language model. We measure improvement on a normalized scale that goes from 0 to 1, with 1 representing perfect overlap between the behavior of the larger pre-trained model and the fine-tuned smaller model, and 0 meaning perfect divergence. Notice that it is in theory possible for $\mathbb{I}$ to go beyond the value of 1, in which case the model is overcorrecting. Conversely, negative values indicate scenarios in which the fine-tuned model exacerbates the behavior of the pre-trained one. It bears pointing out that our measure goes beyond simply observing differences and also factors in the starting similarity of the two pre-trained models: the larger the difference in their behavior, the higher the improvement displayed by the fine-tuned model (if any).
On average, the improvement for the fine-tuned model in within-sample contexts and games amounts to 18%. However, as can be observed in Figure 4, the distribution of improvement in individual scenarios displays considerable heterogeneity. Interestingly, in the case of exacerbated behavior, the fine-tuned model increased its tendency to defect compared to the smaller pre-trained model in two emblematic scenarios. First, it chose to defect when playing Snowdrift under the "friendsharing" context. This seems to suggest that the intrinsically adversarial nature of the game held more weight for it than the friendly social context in which it was taking place. On the contrary, its decision to defect when playing Prisoner’s Delight under the "team" context is harder to parse. Indeed, while the game incentivizes cooperation, it is unclear which logical pathways (if any) lead the model to believe that defection was the more strategically sound option. Arguably, in a reversal of the previous case of exacerbation, it is possible that the model could have perceived the "team" context" as being more adversarial, leading to a decreased rate of cooperation.
We present additional visualizations in SI D.I.
4.2 Out of Sample Results: Contexts
Figure 5: Results for LLaMa2-70b pre-trained, LLaMa2-7b pre-trained, and LLaMa2-7b fine-tuned grouped by game with out-of-sample contexts. We adopt the same conventions as in Fig.3. We keep the structure of the payoffs identical for games both within and out-of-sample. Within each game, and for each context, we observe that the propensity to cooperate with the fine-tuned model always lies in between that of the pre-trained model of the same size and the pre-trained model of the largest size.
Our results for the out-of-sample contexts paint an intricate and complex picture of the effects of fine-tuning. It is immediately clear by way of visual inspection of figure 5 that results are more "well-behaved" than the results for the in-sample contexts, with no trace of overcorrection nor exacerbation. The stark tendency to defect of the larger model when facing the "sports" context is, at a glance, arguably the most surprising result. The model’s consistency in defecting all the time, except when playing Prisoner’s Delight, suggests a hostile understanding of the social frame leading to adversarial behavior. Nevertheless, in line with previous results, LLaMa-70b does shift its behavior when a non-cooperative context is paired with an anti-dilemma.
Figure 6: Improvement of the fine-tuned model over the in-sample games coupled with out-of-sample contexts.
Despite these promising results, Figure 6 shows average improvement is much lower than in the within-sample case. Remarkably, the highest level of improvement is reached when the game is Prisoner’s Delight and the context is "roomsharing". This is possibly because LLaMa2 takes cues from both game structure and context, and in this scenario, the two of them work synergistically to point towards cooperation being the optimal behavior. This theory would also explain why the lowest rate of improvement belongs to the Stag Hunt game played under the "ventcap" scenario: the coordination required to achieve the payoff dominant equilibrium and the social framing of the interaction clash with one another, leading the model to pick the risk-dominant equilibrium.
In conclusion, these nuanced findings demonstrate that the benefits of fine-tuning transcend the specific scenarios presented in the sample. This leads us to believe that smaller models do not merely imitate the data they have been exposed to during training, but also exhibit the ability to generalize to novel, out-of-sample contexts.
4.3 Out of Sample Results: Game and Context
For our final robustness test, we analyze the behavior of the models under consideration in a completely different game structure, with a tailor-made context to accompany it. Experiments conducted on games with different payoffs under within-sample contexts can be found in SI C. In this section, we instead endeavor to understand if the fine-tuned model could generalize its learned behavior beyond the boundaries of two-player games, demonstrating a clearer understanding of social dilemmas and strategic interaction. To accomplish this, we use the Public Good Game (PGG), or the Prisoner’s Dilemma with $n$-players, in order to once again capture the tension between Nash equilibrium and Pareto optimality. Additionally, this allows us to account for scenarios in which choice is not binary, but rather continuous.
Model Mean St. Dev Obs.
LLaMa2-7b 5.49 0.86 300
LLaMa2-7b-FT 5.92 0.99 300
LLaMa2-70b 7 0 300
Table 1: Summary statistics from the Public Pool Game. Mean refers to the mean donation, st. dev. to the standard deviation of mean donation, and obs. is the count of how many instances of the game we run per model.
Results are presented in Table 1. It is immediately evident that LLaMa2-70b exhibits more pro-social behavior and a greater willingness to donate, whereas the 7b pre-trained model displays comparatively selfish tendencies. Notably, the fine-tuned model demonstrates significantly higher levels of pro-social behavior than its pre-trained counterpart, a difference that is statistically significant across all conventional thresholds. This finding suggests that, despite the absence of specific training scenarios involving the game or context, the fine-tuned model retains an enhanced understanding of cooperation, resulting in increased donation levels. Although higher donations may not align with more sophisticated strategic thinking, the friendly interaction environment suggests that the elevated pro-sociality is best attributed to the context-aware intelligence inherent in LLaMa2-70b, a characteristic which has been effectively transferred to the fine-tuned model.
Additional visualizations for all our out-of-sample studies can be found in SI D.II.
5 Conclusion
Throughout the course of this research paper, we have investigated the effectiveness of generating datasets from large pre-trained LLMs for the purpose of instructing smaller models belonging to the same family in tasks that involve theory of mind and strategic thinking. Smaller, strategy-savvy, aligned linguistic models could offer a less costly alternative to consumers without access to powerful graphical processing units without compromising quality of output or performance.
Our research was therefore primarily motivated by our goal of assessing whether smaller models can be fine-tuned to obtain and retain not just the enhanced strategic capabilities of their larger siblings, but also their humanlike attention to social and contextual clues. Concurrently, our attempts to create a dataset to conduct this training find justification in the absence of a canonical corpus of texts that could be used for this purpose. Given LLaMa2-70b’s proven track record as an algorithm capable of balancing strategic considerations with social contexts and clues, we elected to use its responses to a set of 20 unique scenarios as training data. This training data was then used for fine-tuning a smaller algorithm, LLaMa2-7b, by using a memory-efficient technique known as LoRA. In our experiments, the small fine-tuned model departs significantly from the behavior of its pre-trained counterpart and imitates the behavior of a larger pre-trained model of the same family, maintaining considerably high levels of performance even on tasks and when facing contexts that were not present in the training sample.
The significance of these findings is threefold: First, it means that when given the right data in sufficient quantities, a small model can be tuned to mimic the human-aligned strategic behavior of a larger one. Second, it implies that LLMs can themselves be used to generate datasets for ToM tasks, or tasks for which a complete and canonical dataset does not exist. Third, and perhaps most significant, this suggests that fine-tuned models can efficiently and effectively learn to generalize from examples in the training data and display heightened performances even in out-of-sample scenarios and games.
Moreover, this research lays the groundwork for advanced agent-based modeling and simulation techniques, offering new insights into the emergence of pro-social behaviors, such as cooperation [53, 54], trust [55], and fairness [56] in complex systems. By demonstrating the potential of fine-tuned language models as strategic agents, our findings open avenues for creating computationally efficient digital twins of collective human behavior. This approach not only enhances our understanding of cooperation, trust, and fairness in social dynamics but also provides a powerful tool for modeling and analyzing intricate sociotechnical systems. Ultimately, this work bridges the gap between large language model capabilities and practical, scalable simulations of human and organizational behavior [57].
This study represents a promising step towards the development of smaller, aligned, and effective strategic language models. However, our approach does present some limitations. Firstly, the analysis was conducted using a single dataset of fixed size, leaving it unclear whether a larger or smaller dataset would yield different levels of effectiveness. Additionally, while we used a large representative sample of responses from all three models considered, the inherent volatility of high-temperature experiments suggests that a statistical evaluation of strategic choices would benefit from an even larger-scale experiment. Finally, it bears restating that our purpose has been to endow these models with the sophistication necessary to navigate the intersection of social dilemmas and social interactions, yielding results and outcomes that do not always align with what we would expect of a utility-maximizing player. For instance, when playing Prisoner’s Delight, some contexts were associated with an increased rate of defection post-fine-tuning, a choice inconsistent with traditional game theory. In light of this, researchers who would use fine-tuning to generate more sophisticated agents would do well to keep in mind the effect that context and framing inevitably have on these models’ decision-making processes.
In conclusion, this research highlights the need for further exploration into the intersection of language models, game theory, and social interactions. Whether LLMs will prove to be an asset for those wishing to study human behavior will also depend on these algorithms’ capabilities to conceptualize and react to extant structures of incentives which shape and inform interaction in our societies, and the extent to which fine-tuning can empower smaller models to do is yet to be fully understood. The development of small agile language models that can navigate complex social dilemmas has the potential to revolutionize fields such as behavioral economics, decision-making, and social science. As we continue to push the boundaries of what is possible with language models, we may uncover new and innovative ways to study and understand human behavior, and develop more effective and efficient algorithms that can be deployed in a wide range of relevant applications.
Data Availability
The datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request.