### v4

# Large Model Strategic Thinking, Small Model Efficiency: Transferring Theory of Mind in Large Language Models

Abstract
As the performance of larger, newer Large Language Models continues to improve for strategic Theory of Mind (ToM) tasks, the demand for these state-of-the-art models increases commensurately. However, their deployment is costly both in terms of processing power and time. In this paper, we investigate the feasibility of creating smaller, highly-performing specialized algorithms by way of fine-tuning. To do this, we first present a large pre-trained model with 20 unique scenarios that combine different social contexts with games of varying social dilemmas, record its answers, and use them for Q&A fine-tuning on a smaller model of the same family. Our focus is on in-context game-theoretic decision-making, the same domain within which human interaction occurs and that requires both a theory of mind (or a semblance thereof) and an understanding of social dynamics. The smaller model is therefore trained not just on the answers provided, but also on the motivations provided by the larger model, which should contain advice and guidelines to navigate both strategic dilemmas and social cues. We find that the fine-tuned smaller language model consistently bridged the gap in performance between the smaller pre-trained version of the model and its larger relative and that its improvements extended in areas and contexts beyond the ones provided in the training examples, including on out-of-sample scenarios that include completely different game structures. On average for all games, through fine-tuning, the smaller model showed a 46% improvement measured as alignment towards the behavior of the larger model, with 100% representing indistinguishable behavior. When presented with out-of-sample social contexts and games, the fine-tuned model still displays remarkable levels of alignment, reaching an improvement of 18% and 28% respectively. This suggests that our pipeline represents an efficient and intuitive method to transmit some form of theory of mind to smaller models, creating improved and cheaply deployable algorithms in the process. Despite their simplicity and their associated shortcomings and limitations, our findings represent a stepping stone in the pursuit and training of specialized models for strategic and social decision-making.
1 Introduction
The concept of Theory of Mind (ToM) refers to the ability to attribute mental states such as beliefs, intentions, or desires to oneself and others [1]. While extensively studied in humans [2] and other animals [3], ToM has recently gained significant attention in the context of Large Language Models (LLMs) [4, 5, 6, 7]. These sophisticated neural networks, based on Transformers technology, have demonstrated impressive capabilities across various fields, including medicine and finance [8, 9, 10, 11]. The development of ToM capabilities in LLMs is crucial for enabling more natural and efficient human-AI interactions, facilitating the transition to heterogeneous multi-agent LLM systems, and improving LLMs’ ability to understand and respond to complex social contexts. The potential for human-AI interaction and the remarkable performance of LLMs have naturally prompted investigations into their ToM capabilities. However, much like other areas of LLM research, evidence regarding their ToM abilities remains inconclusive, with follow-up studies often challenging or refuting previous findings [12, 13].
While Theory of Mind (ToM) encompasses a wide array of cognitive capabilities, strategic thinking is often highlighted as a particularly specialized and critical aspect of ToM. Indeed, some scholars argue that strategic thinking can serve as a fundamental framework for understanding ToM as a whole [14]. Although larger models do not always yield better outcomes in general ToM tasks [15], they show significant improvements in strategic thinking. This advancement is not merely about enhancing rational, game-theoretic understanding of strategic interactions; it also involves developing nuanced behavior that integrates the stated payoff structure with the context of the game [16, 17]. This integration leads to improved alignment between the strategic behaviors of large language models (LLMs) and humans, which is increasingly important as LLMs are expected to assume a larger role in making consequential strategic decisions, either independently or by assisting human agents [18, 19, 20]. However, despite these promising developments, using large-parameter LLMs for strategic interactions presents challenges. Scaling them up to accommodate a large number of agents, as well as running extensive scenario studies under uncertainty—both critical for practical applications—can become prohibitively expensive, inefficient, and result in substantial environmental implications [21, 22, 23]. Consequently, it would be advantageous to combine the strategic thinking and alignment capabilities of large LLMs with the computational efficiency of smaller models, which ultimately informs the objectives of this paper.
One potential way to address this dual challenge is through fine-tuning. While Large Language Models (LLMs) are inherently generalists, their utility in specialized domains can be limited, as is their demonstration of general and social intelligence. Fine-tuning offers a way to enhance a model’s performance by partially retraining it on targeted datasets or areas of expertise. This approach has shown particular promise in tasks where rote memorization is critical, such as in medical sciences, and could pave the way for creating domain experts "in silico" [24, 25, 26]. However, its effectiveness in open-ended tasks requiring strategic sophistication remains uncertain. In many cases, fine-tuning alone is insufficient to overcome the limitations of even state-of-the-art LLMs in Theory of Mind reasoning [27]. The challenge is further complicated by the lack of suitable datasets for fine-tuning LLMs to enhance strategic behavior without compromising alignment, coupled with the absence of clear guidelines on what such datasets should include or how they should be constructed.
Recent research suggests that larger LLMs can at least navigate simple economic and game theoretic scenarios, [28, 29, 30, 31, 32], but their choices of action tend to be anchored to broader contextual cues [33]. Said cues are occasionally extended by the model itself into fully-fledged hypothetical scenarios, suggesting the importance of context in both human and synthetic decision-making. Contextual cues can be thought of as the motivation for why humans in the real world would make decisions that depart from the predictions at the intersection of game theory and rational choice theory, and the algorithms’ responsiveness to them suggests that these machines might already possess some level of social alignment. In particular, there is evidence that LLaMa2-70b aptly navigates the nuances of the intersection between strategic decision-making and contextual awareness [33].
In light of these findings, we propose a fine-tuning procedure to align the behavior of smaller, more agile algorithms with that of their larger counterpart belonging to the same family. By querying LLaMa2-70b, we obtain a dataset containing motivated responses to four different social dilemmas framed within five different types of social interaction. This dataset is then used to fine-tune LLaMa2-7b, and the performance of the newly fine-tuned model is compared to that of both a pre-trained LLaMa2-70b and a pre-trained LLaMa2-7b. Notably, we find that the smaller fine-tuned model behaves more closely to its larger per-trained counterpart not only for the scenarios upon which it has been trained (i.e. pairwise social dilemma games) but also for new and unknown scenarios and game types that were not present in the training data (e.g. public good game).
These findings highlight that strategic decision-making, a crucial dimension of the Theory of Mind, can be effectively transferred from larger to smaller models. This suggests not only a viable, less computationally demanding alternative to larger models but also the potential for extending this approach to other dimensions of ToM.
2 Literature Review
Model Empowerment and Fine-Tuning
Across a wide range of subjects, the overall consensus is that fine-tuning can dramatically improve the performance of LLMs when queried on specific areas of knowledge [34, 35, 36, 37, 38, 39]. More interestingly, literature is developing on the process of training smaller models using larger models [40, 41, 42], either via fine-tuning or through direct intervention. Our main contribution in this area is in the usage of a completely synthetic dataset for fine-tuning purposes on a domain area that requires strategic thought. Finally, fine-tuning has been shown to improve the performance of large language models even over tasks that require human-like judgment and social skills [43].
LLMs and Strategic Interaction
What characterizes strategic interaction is a necessity to anticipate the moves of a hypothetical co-player - a challenge far different from simply identifying the correct answer to a test. Experiments with LLMs in place of humans tend to yield results that are inconsistent with rational choice theory, behavioral economics, or both [44, 17, 45]. As such, it is currently debated how and where these models should be evaluated [46, 47]. One increasingly popular methodology involves creating LLM-empowered agents and having them compete against one another or against humans, occasionally endowing them with humanlike traits and personality [48, 49, 50]. Unfortunately, this approach is simultaneously too narrow to capture variation in human behavior, and too wide to be generally applicable like game theory is. Nevertheless, there is promise in the usage of these algorithms to mimic human thoughts and behavior [51, 52, 53, 54, 55].
3 Methods
Figure 1: Overview of the methods employed in this paper. We pair all games and scenarios to generate 20 unique combinations, which form the backbone of our dataset. We then submit each combination to each model, and obtain 300 observations per combination. For LLaMa2-70b, we ask for an answer and a motivation; we ask the other models only for their answers. We use the answers coming from LLaMa2-70b to perform LORA on a small, pre-trained LLaMa2-7b. The fine-tuned model is then again queried like the pre-trained model, and once that is done, we collect all data and measure the impact of fine-tuning on preferences.
For the purpose of our analysis, we use models belonging to the LLaMa2 family. Although they no longer represent the state-of-the-art, LLaMa2 models have displayed the least erratic strategic behavior during our investigation and have proven to be much easier to fine-tune. We anticipate that our key findings will apply to more recent models as well. The rationale behind our decision not to use LLaMa3 is discussed in detail in the Supplementary Information (SI) document. (See the SI document, section D). Of the three available configurations, we selected the 70 billion parameters and the 7 billion parameters models to act as the reference large and small pre-trained models, respectively. We loaded all algorithms on [anonymous] University’s High-Performance Cluster (HPC), and we accessed them via the HuggingFace pipeline. We furthermore made use of the langchain package for Python in order to customize system prompts and inputs. Our experiments were conducted on memoryless algorithms with high temperature (0.8), in order to treat each instance and question-answer pair as distinct and to capture the widest possible range in variation.
We distribute a contextual frame of reference upon which to build and elaborate a strategic response to the LLM through the use of system prompts. System prompts serve as an optimal medium for instilling a defined role and personality in LLMs, as they are not processed as typical user inputs but rather function as default settings. This approach is crucial for our objective of grounding LLM decisions within a realistic context. Leveraging system prompts in this manner enhances the ability to align the responses of LLMs with real-world scenarios while not neglecting strategic considerations.
Given that the strategic responses of larger LLMs are shaped by both the stated game payoffs and contextual framing, we adopted the prompting strategy from [33], which demonstrated a wide range of strategic responses across five different contexts for any given payoff matrix. These contexts are: a summit between leaders from two nations ("IR"), a meeting between CEOs of two distinct firms ("biz"), a conference where two industry leaders from different companies commit to environmental regulations ("environment"), a discussion between team members vying for the same promotion ("team"), and a conversation between friends aiming for a compromise ("friendsharing"). To these five "canonical" contexts, we add three more: ("sports"), ("ventcap"), and ("roomsharing"). We call "scenario" the combination of a context and a game; 20 scenarios are used for fine-tuning purposes containing every possible combination of social dilemma and canonical context. 20 more out-of-sample scenarios are generated by combining the five canonical contexts with differently parametrized social dilemmas (with parametrization constant within dilemmas). These pairwise game structures generally fall into four categories: harmony, staghunt, snowdrift, and prisoners’ dilemma [56], representing a spectrum from low to high social dilemma (Our game-theoretic considerations can be found in the SI document, section B.I., game prompts can be found in the SI document, section A.I). Additionally, 12 more out-of-sample scenarios are generated by combining our three additional contexts with the four social dilemmas, whose parametrization is identical to training parametrization. Finally, we run one more scenario which is geared towards multi-player games and which is entirely out-of-sample (both in the game type and the contextual framing). Details of the context prompting are in the SI document, section A.II.
When testing the performance of pre-trained and fine-tuned models on within-sample scenarios, we run 300 initializations and record the response given by the LLM. Prompts are kept constant across LLMs. For LLaMa2-70b, we explicitly ask the algorithm not only to give us its choice of action but also to include a motivation for it. Although the motivations provided by the LLM might include imprecise reasoning and are unlikely to reflect its actual decision-making process, we still chose to include them in the dataset. Given the existing track record on LLaMa2-70b’s behavior, the accompanying elaborations should reflect and corroborate its high social and strategic intelligence with an explanation grounded in natural language. The resulting triplets of system prompt, game prompt, and AI response are then collated within a dataset which is used to fine-tune the smaller LLaMa2-7b model (See the SI document, section B.II, for more details). Additionally, as a robustness check, we ask the larger model to output only its choice of action (without motivation) in order to test for the stability of preferences. Once the fine-tuning has been performed, we query the fine-tuned model both on the scenarios it has been trained on, and on the new out-of-sample ones using 300 initializations per scenario.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: Preliminary investigation of differences in responses between LLaMa2-7b and LLaMa2-70b, grouped by context and game. Clockwise, from the left: propensity to cooperate in LLaMa2-7b grouped by context; propensity to cooperate in LLaMa2-7b grouped by game; propensity to cooperate in LLaMa2-70b grouped by context; propensity to cooperate in LLaMa2-70b grouped by games. Notably, LLaMa2-7b is almost entirely indifferent to context and game and displays a remarkable bias for choosing cooperation, whereas LLaMa2-70b adapts to new contexts and game structures to a remarkable extent.
4 Results
Figure 5 gives an overview of the methods we employ to generate, collect, and then analyze our results.
Figure 3: Improvement (on the $y$ axis) for the fine-tuned version of LLaMa2-7b on within-sample scenarios grouped by (a) game or (b) context. The red vertical line represents an average improvement.
Before initiating the fine-tuning process, we performed an exploratory data analysis to capture the differences in behavior between LLaMa2-7b and LLaMa2-70b. Our results, displayed in figure 2, paint a striking figure: the smallest model in the family is ignorant of both context and game, and outside of its marked preference for choosing cooperation, it seems to act almost randomly. Nowhere is this erratic behavior more visible than in the case of playing Prisoner’s Delight under the friendsharing context. Not only does Prisoner’s Delight admit cooperation from the perspective of a perfectly rational player, but the social environment in which the interaction takes place according to the prompt fed to the machine envisions a scenario in which friends are interacting. In spite of this, LLaMa2-7b chooses to cooperate the least when faced with this combination of game and context when compared to the same game with different contexts. More in general, against all expectations, LLaMa2-7b chooses to cooperate comparatively less when playing under the friendsharing context as opposed to nearly all other contexts. Results obtained from LLaMa2-70b, on the other hand, are perfectly in line with the existing literature, vindicating our choice to use this LLM as a guide through the intricate interplay existing between social context and strategic play.
4.1 Within Sample Results
In order to properly quantify the extent to which the fine-tuned small model approximates the behavior choices of its larger, pre-trained counterpart, we calculate "improvement" as $1-\frac{\mathbb{C}_{70b}-\mathbb{C}_{7b_{Ft}}}{\mathbb{C}_{70b}-\mathbb{C}_{7b}}\times 100$, where $\mathbb{C}_{7b}$ refers to the count of cooperation choices in the original version of LLaMa2-7b model, $\mathbb{C}_{7b_{Ft}}$ refers to the same quantity for the fine-tuned version of the same model, and $\mathbb{C}_{70b}$ captures the count of cooperative choices for the large LLaMa2-70b language model. We measure improvement on a normalized scale that goes from 0 to 1, with 1 representing perfect overlap between the behavior of the larger pre-trained model and the fine-tuned smaller model, and 0 meaning perfect divergence. Notice that it is in theory possible for improvement to go beyond the value of 1, in which case the model is overcorrecting. Conversely, negative values indicate scenarios in which the fine-tuned model exacerbates the behavior of the pre-trained one. It bears pointing out that our measure goes beyond simply observing differences and also factors in the starting similarity of the two pre-trained models: the larger the difference in their behavior, the higher the improvement displayed by the fine-tuned model (if any).
Upon inspection of the results shown in Figure 3, we notice that in aggregate across all games and contexts, the smaller model becomes closer in behavior to the larger model after fine-tuning. On a more granular level, as presented in the SI document, section C.I., we notice that occasionally the fine-tuned model behaves in ways that are not fully consistent with learning. This takes two forms: exacerbation and overcorrection.
Exacerbation occurs when the fine-tuned model amplifies its biases and trends that were present in its pre-trained state in direct opposition to the choices made by the larger teacher model. This is, for instance, the case of playing the Prisoner’s Dilemma under the "team" context, wherein the fine-tuned model cooperates more than its pre-trained counterpart even when the larger model cooperates less. In contrast, overcorrection happens when the fine-tuned model adjusts too strongly in the opposite direction of its initial behavior, to the extent that it surpasses the choices made by the larger model. This, for instance, happens when playing Snowdrift under the "friendsharing" context, with the fine-tuned model cooperating less than the large pre-trained model when the larger pre-trained model would cooperate less than its smaller pre-trained counterpart.
The underlying causes of this erratic behavior remain uncertain and may stem from a variety of factors, including high temperature or inherent model limitations. It is important to note, however, that such inconsistencies are observed in only a subset of the scenarios evaluated. Furthermore, half of the instances of erratic behavior can be attributed to overcorrection, suggesting that the smaller model exhibits a general propensity towards learning, albeit with deviations from expected performance.
4.2 Out-of-Sample Results: Contexts
Figure 4: Improvement for the fine-tuned version of LLaMa2-7b on out-of-sample context and in-sample games grouped by (a) game or (b) context. We adopt the same conventions as in Fig.3. We keep the structure of the payoffs identical for games both within and out-of-sample. As opposed to within-sample scenarios, we observe no exacerbation or overcorrection.
Our exploration of out-of-sample contexts offers a detailed and nuanced understanding of the effects of fine-tuning, revealing both promising trends and inherent challenges. As illustrated in Figure 4 the outcomes in these contexts appear more "well-behaved" compared to the in-sample results, with no evidence of overcorrection or exacerbation. This suggests that fine-tuning contributes to greater stability when the model encounters new data. However, the average improvement observed in these out-of-sample contexts is noticeably lower than in the within-sample cases. This outcome is consistent with our expectations, reflecting a pattern commonly seen in traditional machine learning models where fine-tuned LLaMa2-7b demonstrates solid performance on novel data but falls short of perfection.
Notably, the most pronounced improvements are observed in scenarios featuring anti-dilemmas, such as the Prisoner’s Delight, and in contexts depicting interactions between friends. This finding implies that scenarios characterized by a higher degree of friendliness are more amenable to improvement, with "ventcap", the most antagonistic context, surprisingly ranking third in improvement. On the contrary, the severity of the dilemma itself does not seem to have a straightforward impact on the level of improvement. For instance, despite the Stag Hunt game offering stronger incentives for cooperation than the Prisoner’s Dilemma or Snowdrift, it ranks last in terms of improvement. This suggests a complex interplay between the nature of the dilemma and the effectiveness of fine-tuning, highlighting the need for a more granular understanding of how different contextual factors influence model performance.
In conclusion, these nuanced findings demonstrate that the benefits of fine-tuning transcend the specific scenarios presented in the sample. This leads us to believe that smaller models do not merely imitate the data they have been exposed to during training, but also exhibit the ability to generalize to novel, out-of-sample contexts.
4.3 Out-of-Sample Results: Games
In order to assess whether the fine-tuned model has grasped the underlying dynamics of a certain game, we presented it with out-of-sample games and in-sample contexts. In this case, out-of-sample games are the exact games that the model has already seen in-sample, but with all payoffs scaled by a factor of 2. By scaling the payoffs rather than shifting them, we aim to capture a behavioral response to what is essentially a change in units of measure. Accordingly, we expect a strategically savvy model to keep its responses more or less constant across both in-sample and out-of-sample games. The results of our analysis can be found in Table 1, which reports our statistical tests of significance for observed differences. Notably, in most cases and on average, the fine-tuned model responds to the out-of-sample games in the same way it responded to the in-sample games. Of the 8 scenarios in which a statistically significant difference can be identified, 6 lie at the edge of significance, suggesting that the observed discrepancies might be attributed to noise or high temperature. We conclude that by keeping context fixed, the fine-tuned model can recognize game structures that are fundamentally identical to those already seen in training and act accordingly. This is evidence that fine-tuning can impart a higher level of sophistication on smaller models through the second-order effects of an improved Theory of Mind: the ability to recognize the dynamics underpinning a certain strategic interaction seamlessly extends to interactions that are fundamentally identical save for the unit of measure being used.
Scenario Normal C Ratio OoS C Ratio Difference SE $z$-score $p$-value
team_prison 0.75 0.71 $-0.04$ 0.02 1.6 $0.05^{*}$
team_delight 0.74 0.76 0.01 0.03 $-0.53$ 0.3
team_staghunt 0.74 0.71 $-0.03$ 0.03 1.18 0.12
team_snowdrift 0.70 0.72 0.02 0.03 $-0.88$ 0.19
IR_prison 0.74 0.71 $-0.03$ 0.03 1.05 0.15
IR_delight 0.75 0.70 $-0.05$ 0.02 2.0 $0.02^{*}$
IR_staghunt 0.72 0.69 $-0.03$ 0.03 1.29 0.10
IR_snowdrift 0.71 0.74 0.02 0.03 $-0.89$ 0.19
friendsharing_prison 0.75 0.70 $-0.05$ 0.02 2.0 $0.02^{*}$
friendsharing_delight 0.79 0.74 $-0.05$ 0.02 2.11 $0.02^{*}$
friendsharing_staghunt 0.71 0.77 0.06 0.03 $-2.17$ $0.01^{**}$
friendsharing_snowdrift 0.76 0.76 0.00 0.02 $-0.14$ 0.45
biz_prison 0.74 0.70 $-0.04$ 0.03 1.72 $0.04^{*}$
biz_delight 0.74 0.76 0.03 0.03 $-1.05$ 0.15
biz_staghunt 0.71 0.63 $-0.08$ 0.03 3.06 $0.00^{***}$
biz_snowdrift 0.74 0.76 0.01 0.03 $-0.53$ 0.30
environment_prison 0.63 0.66 0.03 0.03 $-0.96$ 0.17
environment_delight 0.69 0.72 0.03 0.03 $-1.0$ 0.16
environment_staghunt 0.65 0.64 $-0.02$ 0.03 0.61 0.27
environment_snowdrift 0.62 0.67 0.05 0.03 $-1.67$ $0.05^{*}$
AVERAGE 0.72 0.71 $-0.01$ 0.03 0.31 0.38
MEDIAN 0.74 0.71 $-0.01$ 0.03 1.05 0.15
Table 1: Difference in proportion $z$-score testing for propensity to cooperate in within-sample and out-of-sample games for the LLaMa2-7b fine-tuned model. For each scenario, we report: the proportion of cooperative choices in the within-sample game, the proportion of cooperative choices in the out-of-sample game, the difference in proportions, standard error, $z$-score, and associated $p$-value. Reported significance levels follow standard practices: one asterisk (*) for significance at the $0.05$ level, two asterisks (**) for significance at the $0.01$ level, and three asterisks (***) for significance at the $0.001$ level.
4.4 Out-of-Sample Results: Game and Context
Figure 5: Average contribution to the public good for each model, with standard errors in black.
For our final robustness test, we analyze the behavior of the models under consideration in a completely different game structure, with a tailor-made context to accompany it. In this section, we endeavor to understand if the fine-tuned model could generalize its learned behavior beyond the boundaries of two-player games, demonstrating a clearer understanding of social dilemmas and strategic interaction. To accomplish this, we use the Public Good Game (PGG), or the Prisoner’s Dilemma with $n$-players, in order to once again capture the tension between Nash equilibrium and Pareto optimality. Additionally, this allows us to account for scenarios in which choice is not binary, but rather continuous.
Results are presented in Figure 5. It is immediately evident that LLaMa2-70b exhibits more pro-social behavior and a greater willingness to donate, whereas the 7b pre-trained model displays comparatively more selfish tendencies. Notably, the fine-tuned model demonstrates increased levels of pro-social behavior compared to its pre-trained counterpart. This finding suggests that, despite the absence of specific training scenarios involving the game or context, the fine-tuned model retains an enhanced understanding of cooperation, resulting in increased contribution levels. Although higher donations may not align with more sophisticated strategic thinking, the friendly interaction environment suggests that the elevated pro-sociality is best attributed to the context-aware intelligence inherent in LLaMa2-70b, a characteristic that has been effectively transferred to the fine-tuned model. Additional visualizations for all our out-of-sample results can be found in the SI document, section C.II.
5 Conclusion
This research explores the potential of using large pre-trained language models (LLMs) to generate datasets for fine-tuning smaller models in tasks involving theory of mind (ToM) and strategic thinking. Our goal was to assess whether smaller, cost-efficient models could be trained to replicate not just the strategic capabilities of their larger counterparts, but also their nuanced understanding of social and contextual cues, offering a viable alternative to resource-intensive large models.
To address the absence of a canonical dataset for such tasks, we used responses from LLaMa2-70b to 20 unique scenarios as training data. This data was employed to fine-tune a smaller model, LLaMa2-7b, utilizing the state-of-the-art LoRA technique. Our experiments revealed that the fine-tuned smaller model could significantly deviate from its pre-trained behavior, closely imitating the strategic responses of the larger model. Remarkably, the fine-tuned model maintained high performance even in tasks and contexts not present in the training data.
The significance of our findings is threefold: (1) Smaller models, when provided with the right training data, can be fine-tuned to mirror the strategic and social behavior of larger models; (2) LLMs themselves can generate valuable datasets for tasks like ToM, where pre-existing datasets are unavailable; (3) Fine-tuned models exhibit strong generalization capabilities, performing well even in out-of-sample scenarios, which is crucial for their application in real-world settings.
This research paves the way for more efficient LLM-driven agent-based modeling and simulation techniques, which can be used in scenario studies to inform governance and policy decisions in complex sociotechnical systems. By demonstrating the potential of fine-tuned language models as strategic agents, we open up possibilities for creating computationally efficient digital twins of human collective behavior. This approach could significantly enhance our understanding of social dynamics and inform the design of more sophisticated models for studying human and organizational behavior.
Despite the promising results, the study has limitations. The fixed-size dataset used raises questions about the impact of dataset size on model effectiveness, and the variability of high-temperature experiments suggests that larger-scale statistical evaluations are warranted. Moreover, the model’s increased rate of defection in certain contexts, like Prisoner’s Delight, post-fine-tuning, highlights the complex influence of context and framing on decision-making processes. Future researchers should be mindful of these effects when developing sophisticated agents through fine-tuning.
In conclusion, this work underscores the potential of small, strategically aligned LLMs to navigate complex social dilemmas, offering valuable applications in fields like behavioral economics and social science. However, further research is needed to fully understand how fine-tuning can empower smaller models to engage with the incentive structures that drive human interactions. As we advance in this area, we may discover new methodologies for studying and simulating human behavior, leading to the development of efficient, context-aware algorithms for a wide range of applications.
Data Availability
The datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request.